A textbook store sold a combined total of 327 chemistry and physics textbooks in a week. The number of physics textbooks sold was 77 less than the number of chemistry textbooks sold. How many textbooks of each type were sold?
Answer by Cromlix(4381)

Hi,
x = chemistry books
y = physics books
x + y = 327
x - y = 77
Add together
2x = 404
x = 202
Substitute x = 202 into x + y = 327
202 + y = 327
y = 327 - 202
y = 125
Physics books sold = 125
Chemistry books sold = 202
Hope this helps :-)